OpenStudy (anonymous):

Solve the following equation: 7(x-2)+3(x+2)=5(x-6)

OpenStudy (anonymous):

\[7(x-2)+3(x+2)=5(x-6)\] Distribute the 7,3, and 5 \[7x-14+3x+6=5x-30\] Combine like terms \[10x -8=5x-30\] add 8 to both sides. subtract 5x from both sides \[5x = -22\] Divide both sides by 5 \[x = -\frac{ 22 }{ 5 }\]